Obtain the differential equation from the given general solution.

y= c1 cos(4x) +c2 sin(4x) +c3 xcos4x +c4x sin(4x)
1
Expert's answer
2021-09-16T00:49:10-0400

The steps necessary to find the ordinary differential equations satisfied by this solution are


Differentiate the general solution with respect to "x" exactly "4"  times.

Use the "(4+1)" number of expressions ("4" derivatives) obtained to eliminate the "4"  arbitrary constants in terms of the dependent variable or its derivatives.

Obtain the final expression which contains absolutely no arbitrary constant. This is the required differential equation.


"y= c_1 \cos(4x) +c_2 \sin(4x) +c_3 x\cos(4x)"

"+c_4x \sin(4x)"

"y'=-4c_1 \sin(4x) +4c_2 \cos(4x) -4c_3 x\sin4x"

"+c_3 \cos(4x)+4c_4x\cos(4x)+c_4\sin(4x)"

"y''= -16c_1 \cos(4x) -16c_2 \sin(4x) -16c_3 x\cos(4x)"

"-8c_3\sin(4x)-16c_4x\sin(4x)+8c_4\cos(4x)"

"y'''= 64c_1 \sin(4x) -64c_2 \cos(4x)+64c_3 x\sin(4x)"

"-48c_3\cos(4x)-64c_4x\cos(4x)-48c_4\sin(4x)"

"y^{''''}= 256c_1 \cos(4x)+256c_2 \sin(4x) +256c_3 x\cos(4x)"

"+256c_3\sin(4x)+256c_4x\sin(4x)-256c_4\cos(4x)"

"y^{''''}+16y^{''}=128c_3\sin(4x)-128c_4\cos(4x)"

"y^{'''}+16y^{'}=-32c_3\cos(4x)-32c_4\sin(4x)"

"y^{''}+16y= -8c_3\sin(4x)+8c_4\cos(4x)"




"y^{''''}+32y^{''}+256y=0"

The differential equation from the given general solution

"y^{''''}+32y^{''}+256y=0"

Auxiliary (corresponding) equation


"r^2+32r^2+256=0"

"(r^2+16)^2=0"

"r_1=r_2=-4i, r_3=r_4=4i"